OpenStudy (anonymous):

In the circuit below, a constant current of i = 2.09 A flows from b to c through the 8 Ω resistor.

OpenStudy (anonymous):

Calculate the potential difference between points a and d.

OpenStudy (anonymous):

Just learning circuits, I know Ohm's law I=V/R

OpenStudy (anonymous):

FInd the current in the 3 ohms resistor and add it to 2.09 A. The sum will be the total current through the 21 ohms resistor. Now, knowing the current and resistance values you may find the voltage

OpenStudy (anonymous):

Should the current on the 3 ohms resistor be proportional to the resistance?

OpenStudy (anonymous):

What you know is that the voltage over the 3 ohms is the same than voltage over the eifgt ohms. Calculate the current like I=V/R

OpenStudy (anonymous):

What is the current over 3 ohms?

OpenStudy (anonymous):

5.573A

OpenStudy (anonymous):

Assume you have done: I=2.09 x 8 /3

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

good, now what is the total current over the 21 ohms? wouldn't be the sum of both currents?

OpenStudy (anonymous):

so 5.573+ 2.09 = 7.663 A = the current over the 21 resistor

OpenStudy (anonymous):

so now just the total current over total resistance?

OpenStudy (anonymous):

And the voltage over the 21 ohms? wouldn't be 21 x the calculated current?

OpenStudy (anonymous):

yep so voltage over 21= 7.663 * 21= 160.93 V

OpenStudy (anonymous):

And this voltage plus the voltage over the 8 ohms (or over the 3 ohms, that is the same as they are in parallel) wouldn't give you the solution for the voltage between a and d?

OpenStudy (anonymous):

its just V=IR calculate resistance multiply it with I hint resistance in series added simply but in parallel harmonically. . .

OpenStudy (anonymous):

so 160.93V + 2.09A*8ohm= 177.65 V

OpenStudy (anonymous):

If you have done well the operations, then you got it, congratulations!

OpenStudy (anonymous):

Great! that worked. Thank you